1. Convert this equation to linear form T=2 π √ h ^2 +k^2 /gh
1
Expert's answer
2021-02-05T11:47:35-0500

Consider the equation


"T=2\pi\sqrt{\frac{h^2+k^2}{gh}}."

Divide the part inside the root by gh:


"T=\frac{2\pi}{\sqrt{g}}\sqrt{h+\frac{k^2}{h}}."

Multiply the part inside the root by h:


"T=\frac{2\pi}{\sqrt{g}}\sqrt{h^2+k^2}."

If k<<1, we have


"T=\frac{2\pi}{\sqrt{g}}\cdot h."